OpenStudy (travestied):

Describe the transformation that produces the graph of g(x) = 1/2 (x-4)^3 + 5 from the graph of the parent function f(x) = x^3. Give the order in which they must be performed to attain the graph.

OpenStudy (3mar):

you mean \[f(x)=x^3\] has a transformation to be \[g(x)=\frac{ 1 }{ 2 }(x-4)^3+5\]

OpenStudy (3mar):

Let me show you this pic.

OpenStudy (3mar):

\[h(x)=a(bx+c)^3+d\]

OpenStudy (3mar):

b: compress and stretch it horizontally.

OpenStudy (3mar):

C: move it horizontally left and right.

OpenStudy (3mar):

d: move it vertically up and down

OpenStudy (3mar):

a is like b